# Unveiling a 36 Billion Solar Mass Black Hole at the Centre of the Cosmic Horseshoe Gravitational Lens

### v1

@@ 1 Introduction
Most massive galaxies are believed to host a supermassive black hole (SMBH) at their centre. More importantly, host galaxies and their SMBHs exhibit clear scaling relations, pointing to a co-evolution between the galaxy and the black hole (BH; Kormendy & Ho, 2013). The BH mass ($M_{\text{BH}}$) has been shown to correlate with various galaxy properties, such as the bulge luminosity (e.g., Magorrian et al., 1998; Marconi & Hunt, 2003; Gültekin et al., 2009), stellar bulge mass (e.g., Laor, 2001; McLure & Dunlop, 2002), dark matter (DM) halo mass (e.g., Marasco et al., 2021; Powell et al., 2022), and stellar velocity dispersion (e.g., Gebhardt et al., 2000; Beifiori et al., 2009). Notably, the $M_{\text{BH}}-\sigma_{e}$ relation, which links BH mass to the effective stellar velocity dispersion of the host ($\sigma_{e}$), remains tight across various morphological types and BH masses (van den Bosch, 2016). Nonetheless, when SMBHs accrete mass from their neighbourhoods, they can act as active galactic nuclei (AGNs), injecting energy in the surrounding gas in a form of feedback. This feedback can be either positive, triggering star formation (Ishibashi & Fabian, 2012; Silk, 2013), or negative, quenching galaxy growth (e.g., Hopkins et al., 2006; Dubois et al., 2013).
@@
It is expected that the most massive galaxies in the Universe, such as brightest cluster galaxies (BCGs), host the most massive SMBHs. Indeed, so-called ultramassive black holes (UMBHs; $M_{\text{BH}}\geq 10^{10}M_{\odot}$) have been found in such systems (e.g., Hlavacek-Larrondo et al., 2012). Most of these UMBHs have been measured through spatially resolved dynamical modelling of stars and/or gas. For instance, the UMBH in Holm 15A at $z=0.055$ ($M_{\text{BH}}=(4.0\pm 0.80)\times 10^{10}M_{\odot}$; Mehrgan et al., 2019) and the UMBH in NGC 4889 at $z=0.021$ ($M_{\text{BH}}=(2.1\pm 1.6)\times 10^{10}M_{\odot}$; McConnell et al., 2012) were both determined using stellar dynamical modelling. However, despite the success of this technique in yielding hundreds of BH mass measurements, the requirement for high-quality spatially resolved spectroscopy poses significant challenges for studies at increasing redshift (see, e.g., Kormendy & Ho, 2013, Suplemental Material S1).
@@
Nonetheless, the significance of these UMBHs lies in the fact that many of them deviate from the standard linear $M_{\text{BH}}-\sigma_{e}$ relation (e.g., Kormendy & Ho, 2013; van den Bosch, 2016). This suggests either a distinct evolutionary mechanism governing the growth of the largest galaxies and their BHs (McConnell et al., 2011), leading to a significantly steeper relation (Bogdán et al., 2018), or a potential decoupling between the SMBH and host galaxy co-evolution. Populating the high-mass end of the $M_{\text{BH}}-\sigma_{e}$ relation, particularly through direct BH measurements, could help resolve this ongoing puzzle.
@@
Recently, Nightingale et al. (2023), by modelling the gravitationally lensed radial image in the Abell 1201 BCG ($z=0.169$), was able to measure the mass of its dormant SMBH as $M_{\text{BH}}=(3.27\pm 2.12)\times 10^{10}M_{\odot}$. This provides a complementary approach to other higher-z probes of BH mass, such as reverberation mapping (Blandford & McKee, 1982; Bentz & Katz, 2015) and AGN spectral fitting (Shen, 2013). Unlike these methods, which require active accretion and depend on local Universe calibrations, the lensing technique offers a direct measurement independent of the SMBH’s accretion state.
@@
In this paper, we analyse the Cosmic Horseshoe gravitational lens system (Belokurov et al., 2007), where the lens galaxy is one of the most massive strong gravitational lenses known. The lens galaxy is an early-type galaxy (ETG) at redshift $z_{l}=0.44$, possibly part of a fossil group, and is notable for lensing one of its sources into a nearly complete Einstein ring (the Horseshoe). Additionally, a second multiply imaged source forms a radial arc near the centre of the lens galaxy.
@@
Thanks to the radial image formed very close to the centre, the inner DM distribution of the Cosmic Horseshoe can be studied in detail, as demonstrated by Schuldt et al. (2019). By simultaneously modelling stellar kinematics from long-slit spectroscopy and the positions of the lensed sources, Schuldt et al. (2019) found that the DM halo is consistent with a Navarro-Frenk-White (NFW; Navarro et al., 1997) profile, with the DM fraction within the Einstein radius ($R_{\text{Ein}}$) estimated to be between $60\%$ and $70\%$. Moreover, their models include a point mass at the galaxy’s centre, reaching values around $\sim 10^{10}M_{\odot}$, which could represent a SMBH; however, they did not pursue further investigations into this possibility.
@@
Using new integral-field spectroscopic data from the Multi Unit Spectroscopic Explorer (MUSE) and imaging from the Hubble Space Telescope (HST), we conducted a systematic modelling of the Cosmic Horseshoe system to reassess the evidence for an SMBH at the heart of the lens galaxy. We performed a self-consistent analysis of both strong gravitational lensing (SGL) and stellar dynamics, which demonstrated that the presence of an UMBH is necessary to fit both datasets simultaneously. Additionally, we found that including the radial arc improves the accuracy of the BH mass estimation, compared to models based solely on stellar dynamics.
@@
This paper is structured as follows: In Section 2, we present the HST imaging data and MUSE observations, along with the kinematic maps used for the dynamical modelling. Section 3 briefly summarises the lensing and dynamical modelling techniques, including the multiple-lens-plane formalism, the approximations adopted in this work, and the mass profile parametrisation. In Section 4, we present the results from our fiducial model and alternatives models, which we use to address the systematics on the BH mass. In Section 5 we discuss our results and present other astrophysical implications. Finally, we summarise and conclude in Section 6.
Throughout this paper, we adopt the cosmological parameters consistent with Planck Collaboration et al. (2016): $\Omega_{\Lambda,0}=0.6911$, $\Omega_{\text{m},0}=0.3089$, $\Omega_{\text{b},0}=0.0486$, and $H_{0}=67.74$ km s-1 Mpc-1.
@@ 2 Data
The Cosmic Horseshoe (SDSS J1148+1930) was first identified by Belokurov et al. (2007) as part of the Sloan Digital Sky Survey (SDSS). The primary lens is a massive ETG at $z_{l}=0.44$, with an estimated enclosed mass within $R_{\text{Ein}}$ of $\sim 5\times 10^{12}M_{\odot}$ (Dye et al., 2008; Schuldt et al., 2019). The radial arc and its counter-image correspond to a source at redshift $z_{s1}=1.961$ (s1, hereafter), while the tangential arc is a star-forming galaxy (James et al., 2018) at redshift $z_{s2}=2.381$ (s2, hereafter). In Fig. 1, the radial arc and its counter-image are highlighted within white boxes, with a zoomed-in view of the radial image displayed in the inset.
Figure 1: HST/WFC3 colour composite image of the Cosmic Horseshoe, created using the F814W, F606W, and F475W filters. The system is composed by the main deflector ($z_{l}=0.44$); the eponymous Einstein ring of the Cosmic Horseshoe ($z_{\text{s2}}=2.381$); and the radial arc and its counter-image ($z_{s1}=1.961$), both highlighted. The inset shows the radial arc. It is worth noting that, even in the F475W filter where s1 is brighter, the radial arc remains very faint. The figure is oriented such that east is up and north is right.
@@ 2.1 HST imaging
To construct the point spread function (PSF) for each filter, we identified non-saturated stars from the Gaia DR2 catalogue (Arenou et al., 2018) and performed interactive PSF modelling using PSFr (see, e.g., Birrer et al., 2019). As in Schuldt et al. (2019), the noise map for each pixel $i$ was calculated by combining the background noise $\sigma_{\text{bkgd}}$ and Poisson noise in quadrature: $\sigma^{2}_{\text{rms,i}}=\sigma^{2}_{\text{bkgd}}+\sigma^{2}_{\text{Poisson,i}}$. The background level was estimated as a constant value, measured from an empty region of sky near the main deflector, using the astropy sigma-clipping method. The Poisson noise was calculated from the effective exposure map and the intensity counts for each pixel.
@@ 2.2 MUSE observations and kinematical map
The mean velocity ($v$) and velocity dispersion ($\sigma_{v}$) in each Voronoi bin were measured using the penalized pixel fitting method, as implemented in the pPXF software (Cappellari & Emsellem, 2004). For templates, we adopted the X-shooter Spectral Library SSP models (XLS-SSP; Verro et al., 2022), which offer a resolution of $\sim 13$ km s-1 and a wavelength coverage from $350$ to $2480$ nm. The XLS-SSP library was selected due to its high resolution, enabling convolution with the MUSE instrumental resolution after de-shifting the galaxy spectra. Each Voronoi bin was fitted over the wavelength range $5600-7600$  (galaxy-frame), with emission lines within this range masked. Uncertainties for $v$ and $\sigma_{v}$ were determined through Monte Carlo perturbations of the best-fit model, performing $200$ realisations and taking the standard deviation as the uncertainty.
@@ 3 Methods
We construct a fully self-consistent mass model for the main lensing galaxy by jointly modelling its SGL effect and spatially resolved stellar velocity dispersion. This combined approach has been successfully applied in previous studies to constrain the mass profiles of ETGSs in tests of modified gravity (e.g., Collett et al., 2018; Melo-Carneiro et al., 2023) and to investigate the distribution of baryonic and DM content within galaxies (e.g., Barnabè et al., 2012; Wang et al., 2022). By simultaneously leveraging both methods, we mitigate degeneracies that arise when using them independently, such as the mass-anisotropy degeneracy (e.g., Gerhard, 1993) and the mass-sheet degeneracy (e.g., Gorenstein et al., 1988).
@@ 3.1.1 Formalism
Here, $\kappa=\Sigma/\Sigma_{\text{crit}}$ represents the surface mass density of the lens scaled by the critical surface density, $\Sigma_{\text{crit}}$, defined as:
@@
where $D_{ij}$ is the angular diameter distance between observer, lens, and source.
@@
${\theta}_{j}=\mn@boldsymbol{\theta}_{1}-\sum_{i=1}^{j-1}\eta_{ij}\mn@boldsymbol{\alpha}_{i}\left(\mn@boldsymbol{\theta}_{i}\right),$ (5)
where ${\theta}_{j}$ is the angular position of a light ray in the $j^{\text{th}}$ plane, and ${\alpha}_{i}$ is the angular deflection caused by the $i^{\text{th}}$ lens acting on rays originating from the furthest redshift source plane. The factor $\eta_{ij}$ is the scaling factor as defined in Eq. 4, i.e.,
$\eta_{ij}=\frac{D_{ij}}{D_{j}}\frac{D_{s}}{D_{is}},$ (6)
with $s$ being the most distant source. In the case of just one lens and one source, Eq. 5 reduces to Eq. 1, with ${\theta}_{2}=\mn@boldsymbol{\beta}$, ${\theta}_{1}=\mn@boldsymbol{\theta}$, and $\eta_{ij}=1$.
To reconstruct the unlensed sources in practice, the lens model should incorporate the deflections caused by s1 on s2. However, for the Cosmic Horseshoe system, the deflection contribution from s1 to the total angle experienced by s2 is expected to be negligible for three primary reasons: (i) the sources are at similar redshifts, so s1 will be an inefficient lens regardless of its mass (ii) The location of the radial image and its counter-image, imply that s1 is not closely aligned with s2 and that the images of s2 do not pass close to s1; and (iii) the low observed lensed surface brightness of s1, combined with its small size after reconstruction (see Section 4), suggests that it is a low-mass galaxy. Given these assumptions, in this work, we neglect the lensing effect of s1 on s2.
@@ 3.1.2 Lens modelling
For the lens modelling, we employ the open-source software PyAutoLens (Nightingale et al., 2018, 2021), which implements a Bayesian version (Suyu et al., 2006) of the semi-linear inversion (SLI) method (Warren & Dye, 2003). For a given set of non-linear parameters (describing the lens mass model and/or the source), the code linearly ray-traces image-pixels from the image plane back to the source plane, reconstructing the source emission using an adaptive mesh grid [^8] This assumes that the lens light was subtracted already.
@@
In PyAutoLens the native likelihood is that described in Suyu et al. (2006, eq. 19) and Dye et al. (2008, eq. 5). This approach incorporates the instrumental PSF blurring alongside regularisation terms for the pixelised source reconstruction, which helps to mitigate ill-posed solutions during the linear inversion. We refer the readers to these references for more details.  [^9] Those interested in the technical aspects and implementation within PyAutoLens may also check the following notebooks: https://github.com/Jammy2211/autolens_workspace/tree/main/notebooks/advanced/log_likelihood_function.
@@
To remove over- and under-magnified solutions (Maresca et al., 2021), we trace two image plane regions of s1 to the source plane, where they are expected to overlap. If these conjugate regions fail to overlap after delensing, we penalise the likelihood of this solution by a factor proportional to the distance between the two regions in the source plane. This approach effectively removes unphysical solutions resembling over/under-magnified versions of the data, without the need for parameter fine-tuning. Using regions, rather than pairs of conjugate points, also mitigates the risk of selecting incorrect pairs. This is particularly relevant in the case of the Horseshoe system, where the radial image is very faint and embedded within the lens light, making it challenging to reliably identify conjugate pairs between the radial image and its counter-image. Further details of this approach are provided in Appendix A.
In this work, we did not attempt to model the full Einstein ring of s2, as our primary focus is the inner mass distribution of the main deflector, which is better constrained by the lensed radial image of s1. However, the Einstein ring provides valuable information about the large-scale mass distribution, offering an additional constraint on the mass profile. Furthermore, the Einstein ring is sensitive to the mass of s1, as it reflects the total projected mass distribution along the LOS between the main-lens and s2.
@@
To incorporate the lensing information from s2, we apply a Gaussian prior on the total mass enclosed within its Einstein radius, $M_{\text{Ein}}$. This prior has a mean of $M_{\text{Ein}}=5.46\times 10^{12}M_{\sun}$, corresponding to an effective Einstein radius of $R_{\text{Ein}}=5.08$″, and a standard deviation of $\sigma_{M_{\text{Ein}}}=0.27\times 10^{12}M_{\sun}$. We derived $M_{\text{Ein}}$ by modelling the lensed s2 source using an elliptical power-law (EPL) mass profile for the main deflector with PyAutoLens. The resulting values for $M_{\text{Ein}}$ and $R_{\text{Ein}}$ are consistent with those reported by Dye et al. (2008) and Schuldt et al. (2019). Fig. 3 shows the data pixels used to model the Cosmic Horseshoe, along with the highest-likelihood image model and the reconstructed source. Further details about the EPL mass model are provided in Appendix B.
It is important to highlight that the $M_{\text{Ein}}$ prior accounts for the mass perturbation from s1. Specifically, $M_{\text{Ein}}$ represents the total mass within a cylinder of radius $R_{\text{Ein}}$, centred on the main-lens, along the LOS between the observer and s2. Nevertheless, as discussed in Section 4, excluding this prior from the likelihood does not impact the inferred BH mass.

### v3

@@ 1 Introduction
Most massive galaxies are believed to host a supermassive black hole (SMBH) at their centre. More importantly, host galaxies and their SMBHs exhibit clear scaling relations, pointing to a co-evolution between the galaxy and the SMBH (Kormendy & Ho, 2013). The SMBH mass ($M_{\text{BH}}$) has been shown to correlate with various galaxy properties, such as the bulge luminosity (e.g., Magorrian et al., 1998; Marconi & Hunt, 2003; Gültekin et al., 2009), stellar bulge mass (e.g., Laor, 2001; McLure & Dunlop, 2002), dark matter (DM) halo mass (e.g., Marasco et al., 2021; Powell et al., 2022), number of host’s globular clusters (e.g., Burkert & Tremaine, 2010; Harris et al., 2014), and stellar velocity dispersion (e.g., Gebhardt et al., 2000; Beifiori et al., 2009). Notably, the $M_{\text{BH}}-\sigma_{e}$ relation, which links SMBH mass to the effective stellar velocity dispersion of the host ($\sigma_{e}$), remains tight across various morphological types and SMBH masses (van den Bosch, 2016). Nonetheless, when SMBHs accrete mass from their neighbourhoods, they can act as active galactic nuclei (AGNs), injecting energy in the surrounding gas in a form of feedback. This feedback can be either positive, triggering star formation (Ishibashi & Fabian, 2012; Silk, 2013; Riffel et al., 2024), or negative, quenching galaxy growth (e.g., Hopkins et al., 2006; Dubois et al., 2013; Costa-Souza et al., 2024).
@@
It is expected that the most massive galaxies in the Universe, such as brightest cluster galaxies (BCGs), host the most massive SMBHs. Indeed, so-called ultramassive black holes (UMBHs; $M_{\text{BH}}\geq 10^{10}M_{\odot}$) have been found in such systems (e.g., Hlavacek-Larrondo et al., 2012). Most of these UMBHs have been measured through spatially-resolved dynamical modelling of stars and/or gas. For instance, the UMBH in Holm 15A at $z=0.055$ ($M_{\text{BH}}=(4.0\pm 0.80)\times 10^{10}M_{\odot}$; Mehrgan et al., 2019) and the UMBH in NGC 4889 at $z=0.021$ ($M_{\text{BH}}=(2.1\pm 1.6)\times 10^{10}M_{\odot}$; McConnell et al., 2012) were both determined using stellar dynamical modelling. However, despite the success of this technique in yielding hundreds of SMBH mass measurements, the requirement for high-quality spatially resolved spectroscopy poses significant challenges for studies at increasing redshift (see, e.g., Kormendy & Ho, 2013, Suplemental Material S1).
@@
Nonetheless, the significance of these UMBHs lies in the fact that many of them deviate from the standard linear $M_{\text{BH}}-\sigma_{e}$ relation (e.g., Kormendy & Ho, 2013; van den Bosch, 2016). This suggests either a distinct evolutionary mechanism governing the growth of the largest galaxies and their SMBHs (McConnell et al., 2011), leading to a significantly steeper relation (Bogdán et al., 2018), or a potential decoupling between the SMBH and host galaxy co-evolution. Populating the high-mass end of the $M_{\text{BH}}-\sigma_{e}$ relation, particularly through direct $M_{\text{BH}}$ measurements, could help resolve this ongoing puzzle.
@@
Recently, Nightingale et al. (2023), by modelling the gravitationally lensed radial image near the the Abell 1201 BCG ($z=0.169$), was able to measure the mass of its dormant SMBH as $M_{\text{BH}}=(3.27\pm 2.12)\times 10^{10}M_{\odot}$, therefore an UMBH. This provides a complementary approach to other high-$z$ probes of SMBH mass, such as reverberation mapping (Blandford & McKee, 1982; Bentz & Katz, 2015) and AGN spectral fitting (Shen, 2013). Unlike these methods, which require active accretion and depend on local Universe calibrations, the lensing technique offers a direct measurement independent of the SMBH’s accretion state.
@@
In this paper, we analyse the Cosmic Horseshoe gravitational lens system (Belokurov et al., 2007), where the lens galaxy is one of the most massive strong gravitational lenses known to date. The lens galaxy is an early-type galaxy (ETG) at redshift $z_{l}=0.44$, possibly part of a fossil group (Ponman et al., 1994), and is notable for lensing one of its sources into a nearly complete Einstein ring (the Horseshoe). Additionally, a second multiply imaged source forms a radial arc near the centre of the lens galaxy.
@@
Thanks to the radial image formed very close to the centre, the inner DM distribution of the Cosmic Horseshoe can be studied in detail, as done by Schuldt et al. (2019). By simultaneously modelling stellar kinematics from long-slit spectroscopy and the positions of the lensed sources, Schuldt et al. (2019) found that the DM halo is consistent with a Navarro-Frenk-White (NFW; Navarro et al., 1997) profile, with the DM fraction within the effective radius ($R_{e}$) estimated to be between $60\%$ and $70\%$. Moreover, their models include a point mass at the galaxy’s centre, reaching values around $\sim 10^{10}M_{\odot}$, which could represent an SMBH; however, they did not pursue further investigations into this possibility.
@@
Using new integral-field spectroscopic data from the Multi Unit Spectroscopic Explorer (MUSE) and imaging from the Hubble Space Telescope (HST), we conducted a systematic modelling of the Cosmic Horseshoe system to reassess the evidence for an SMBH at the heart of the lens galaxy. We performed a self-consistent analysis of both strong gravitational lensing (SGL) and stellar dynamics, which demonstrated that the presence of an SMBH is necessary to fit both datasets simultaneously.
@@
This paper is structured as follows: In Section 2, we present the HST imaging data and MUSE observations, along with the kinematic maps used for the dynamical modelling. Section 3 briefly summarises the lensing and dynamical modelling techniques, including the multiple-lens-plane formalism, the approximations adopted in this work, and the mass profile parametrisation. In Section 4, we present the results from our fiducial model and alternatives models, which we use to address the systematics on the SMBH mass. In Section 5 we discuss our results and present other astrophysical implications. Finally, we summarise and conclude in Section 6.
Unless otherwise, all parameter estimates are derived from the final sampling chain, with reported values representing the median of each parameter’s one-dimensional marginalised posterior distribution, with uncertainties corresponding to the $16^{\text{th}}$ and $84^{\text{th}}$ percentiles. Furthermore, throughout this paper, we adopt the cosmological parameters consistent with Planck Collaboration et al. (2016): $\Omega_{\Lambda,0}=0.6911$, $\Omega_{\text{m},0}=0.3089$, $\Omega_{\text{b},0}=0.0486$, and $H_{0}=67.74$ km s-1 Mpc-1.
@@ 2 Data
The Cosmic Horseshoe (SDSS J1148+1930) system was first identified by Belokurov et al. (2007) as part of the Sloan Digital Sky Survey (SDSS). The primary lens is a massive ETG at $z_{l}=0.44$, with an estimated enclosed mass within the Einstein radius ($R_{\text{Ein}}$) of $\sim 5\times 10^{12}M_{\odot}$ (Dye et al., 2008; Schuldt et al., 2019). The radial arc and its counter-image correspond to a source at redshift $z_{s1}=1.961$ (s1, hereafter), while the tangential arc is a star-forming galaxy (James et al., 2018) at redshift $z_{s2}=2.381$ (s2, hereafter). In Fig. 1, the radial arc and its counter-image are highlighted within white boxes, with a zoomed-in view of the radial image displayed in the inset.
Figure 1: HST/WFC3 colour composite image of the Cosmic Horseshoe, created using the F814W, F606W, and F475W filters. The system is composed by the main deflector ($z_{l}=0.44$); the eponymous Einstein ring of the Cosmic Horseshoe ($z_{\text{s2}}=2.381$); and the radial arc and its counter-image ($z_{s1}=1.961$), both highlighted. The inset shows the radial arc. The figure is oriented such that north is up and east is left.
@@ 2.1 HST imaging
To construct the point spread function (PSF) for each filter, we identified two non-saturated stars from the Gaia DR2 catalogue (Arenou et al., 2018) and performed a interactive PSF modelling using PSFr (see, e.g., Birrer et al., 2019). We follow the same procedure as Birrer et al. (2019), applying a $90^{\circ}$ rotation symmetry based on the HST optical symmetry. This allows each star to be rotated four times, enabling PSF model estimation using a total of eight stacked images. The noise map for each pixel $i$ was calculated by combining the background noise $\sigma_{\text{bkgd}}$ and Poisson noise in quadrature: $\sigma^{2}_{\text{rms,i}}=\sigma^{2}_{\text{bkgd}}+\sigma^{2}_{\text{Poisson,i}}$. The background level was estimated as a constant value, measured from an empty region of sky near the main deflector, using the astropy sigma-clipping method. The Poisson noise was calculated from the effective exposure map and the intensity counts for each pixel.
@@ 2.2 MUSE observations and kinematical map
The mean velocity ($v$) and velocity dispersion ($\sigma_{v}$) in each Voronoi bin were measured using the penalized pixel fitting method, as implemented in the pPXF software (Cappellari & Emsellem, 2004). For templates, we adopted the full X-shooter Spectral Library SSP models (XLS-SSP-DR3; Verro et al., 2022), which offer a resolution of $\sim 13$ km s-1 and a wavelength coverage from $3500$ to $24800$. The XLS-SSP library was selected due to its high resolution, enabling convolution with the MUSE instrumental resolution after de-shifting the galaxy spectra. Each Voronoi bin was fitted over the wavelength range $5600-7600$  (galaxy-frame), with emission lines within this range masked.
Uncertainties for $v$ and $\sigma_{v}$ were determined through Monte Carlo perturbations of the best-fit model, performing $200$ realisations and taking the standard deviation as the uncertainty. Template selection is a known source of systematic uncertainty. Recently, Knabel et al. (2025) highlighted that template mismatch can dominate the error budget in high-precision applications, for instance in time-delay cosmography. However, for SMBH mass measurements, we consider its impact to be secondary relative to other sources of uncertainty, such as assumptions in the dynamical modelling (e.g., anisotropy profile and intrinsic shape). A detailed investigation of template choice effects on SMBH mass measurements is beyond the scope of this work.
@@ 3 Methods
We construct a fully self-consistent mass model for the main lensing galaxy by jointly modelling its SGL effect and spatially-resolved stellar velocity dispersion. This combined approach has been successfully applied in previous studies to constrain the mass profiles of ETGs in tests of modified gravity (e.g., Collett et al., 2018; Melo-Carneiro et al., 2023) and to investigate the distribution of baryonic and DM content within galaxies (e.g., Barnabè et al., 2012; Wang et al., 2022). By simultaneously leveraging both methods, we mitigate degeneracies that arise when using them independently, such as the mass-anisotropy degeneracy (e.g., Gerhard, 1993) and the mass-sheet degeneracy (e.g., Gorenstein et al., 1988). Furthermore, the Cosmic Horseshoe system is a compound lens with two sources at different redshifts (Schuldt et al., 2019). The joint modeling of these sources can also help break (or at least alleviate) these degeneracies present in the lens model (e.g., Enzi et al., 2025).
@@ 3.1.1 Formalism
Here, the convergence $\kappa=\Sigma/\Sigma_{\text{crit}}$, represents the surface mass density of the lens scaled by the critical surface density, $\Sigma_{\text{crit}}$, defined as:
@@
where $D_{ls}$, $D_{l}$ and $D_{s}$ are the angular diameter distances between the lens and source, lens and observer, and source and observer, respectively.
@@
${\theta}_{k}=\mn@boldsymbol{\theta}_{1}-\sum_{j=1}^{k-1}\eta_{jk}\mn@boldsymbol{\alpha}_{j}\left(\mn@boldsymbol{\theta}_{j}\right),$ (5)
where ${\theta}_{k}$ is the angular position of a light ray in the $k^{\text{th}}$ plane, and ${\alpha}_{j}$ is the angular deflection caused by the $j^{\text{th}}$ lens acting on rays originating from the furthest redshift source plane. The factor $\eta_{jk}$ is the scaling factor as defined in Eq. 4, i.e.,
$\eta_{jk}=\frac{D_{jk}}{D_{k}}\frac{D_{s}}{D_{js}},$ (6)
with $s$ being the most distant source. In the case of just one lens and one source, Eq. 5 reduces to Eq. 1, with ${\theta}_{2}=\mn@boldsymbol{\beta}$, ${\theta}_{1}=\mn@boldsymbol{\theta}$, and $\eta_{jk}=1$.
To reconstruct the unlensed sources, the lens model must, in principle, account for deflections from both the main deflector and source s1 on s2. However, in the Cosmic Horseshoe system, the deflection contribution from s1 to the total deflection angle experienced by s2 is expected to be negligible for three main reasons: (i) the sources lie at similar redshifts, making s1 an inefficient lens; (ii) the positions of the radial image and its counter-image indicates that s1 is not closely aligned with s2, and the lensed images of s2 do not pass near s1; and (iii) the low observed lensed surface brightness of s1, combined with its small size after reconstruction (see Section 4), suggests that it is a low-mass galaxy. To further support this assumption, we perform a compound lens model including s1’s mass contribution, which confirms its negligible impact.
@@ 3.1.2 Lens modelling
For the lens modelling, we employ the open-source software PyAutoLens (Nightingale et al., 2018, 2021), which implements a Bayesian version (Suyu et al., 2006) of the semi-linear inversion (SLI) method (Warren & Dye, 2003). For a given set of non-linear parameters (describing the lens mass model and/or the source), the code linearly ray-traces image-pixels from the image plane back to the source plane, reconstructing the source emission using an adaptive mesh grid [^8] This assumes that the lens light was subtracted already..
@@
In PyAutoLens the native likelihood is that described in Suyu et al. (2006, eq. 19) and Dye et al. (2008, eq. 5). This approach incorporates the instrumental PSF blurring alongside regularisation terms for the pixelised source reconstruction, which helps to mitigate ill-posed solutions during the linear inversion. We refer the readers to these references for more details  [^9] Those interested in the technical aspects and implementation within PyAutoLens may also check the following notebooks: https://github.com/Jammy2211/autolens_workspace/tree/main/notebooks/advanced/log_likelihood_function.
@@
To remove over- and under-magnified solutions (Maresca et al., 2021), we trace two image plane regions of s1 back to the source plane, where they are expected to overlap. If these conjugate regions fail to overlap after delensing, we penalise the likelihood of this solution by a factor proportional to the distance between the two regions in the source plane. This approach effectively removes unphysical solutions resembling over/under-magnified versions of the data, without the need for parameter fine-tuning. Using regions, rather than pairs of conjugate points, also mitigates the risk of selecting incorrect pairs. This is particularly relevant in the case of the Cosmic Horseshoe system, where the radial image is very faint and embedded within the lens light, making it challenging to reliably identify conjugate pairs between the radial image and its counter-image. Further details of this approach are provided in Appendix A.
In this work, our primary goal is to constrain the inner mass distribution of the main deflector through joint modelling of the radial image of s1 and its counter-image, along with the host galaxy’s stellar kinematics. As a result, we do not attempt to model the full Einstein ring of s2 in our main analysis. Nonetheless, the lensed emission from s2 traces the total projected mass distribution on larger scales and thus provides valuable complementary constraints on the global mass profile of the main deflector.
@@
To incorporate this information, we first modeled the lensed emission from source s2 independently, using an elliptical power-law (EPL) mass profile for the main deflector implemented in PyAutoLens. This fit yields a mass within the Einstein radius of $M_{\text{Ein}}=5.46\times 10^{12}\,M_{\sun}$, corresponding to an effective Einstein radius of $R_{\text{Ein}}=5.08\arcsec$. These values agree with those reported by Dye et al. (2008) and Schuldt et al. (2019). We then incorporate this result into our combined lensing+dynamics analysis by imposing a Gaussian prior on $M_{\text{Ein}}$ with mean $5.46\times 10^{12}M_{\sun}$ and standard deviation of $\sigma_{M_{\text{Ein}}}=0.27\times 10^{12}M_{\sun}$. Figure 3 shows the image pixels used for the EPL fit, the highest-likelihood model, and the reconstructed source. Further details of the EPL fitting procedure are provided in Appendix B.
To test the assumption that compound lensing by the first source is negligible, we also ran a compound lens model with the HercuLens code (Galan et al., 2022; Enzi et al., 2025), including a singular isothermal ellipse mass component for s1. For this compound model the Einstein radius of the primary lens never changed by more than 0.6%, which is substantially smaller than the width of the Gaussian we have adopted for the mass of the primary lens within the main Horseshoe.
While this prior provides an additional constraint on the large-scale mass distribution, we test its impact in Section 4 and find that excluding it from the likelihood has no effect on the inferred SMBH mass.